# Identification and Control of Neutral Anyons

### v1

@@
Identification and Control of Neutral Anyons
@@
Ron Q. Nguyen1    Naiyuan J. Zhang1    Navketan Batra1,2    Xiaoxue Liu3    Kenji Watanabe4    Takashi Taniguchi5    D. E. Feldman1,2    J.I.A. Li1 1Department of Physics, Brown University, Providence, Rhode Island 02912, USA 2Brown Theoretical Physics Center, Brown University, Providence, Rhode Island 02912, USA 3Tsung-Dao Lee Institute and School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China 4Research Center for Electronic and Optical Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 5Research Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(October 31, 2024)
Beyond the well-known fermions and bosons, anyons—an exotic class of particles—emerge in the fractional quantum Hall effect and exhibit fractional quantum statistics. Anyons can be categorized by their charge, with extensive research focused on those carrying fractional charge, while charge-neutral anyons in 2D electron liquids remain largely unexplored. Here, we introduce bilayer excitons as a new pathway to realizing charge-neutral anyons. By pairing quasiparticles and quasiholes from Laughlin states, we report bilayer excitons that obey fractional quantum statistics. Through layer-asymmetric field-effect doping, we achieve precise control of the anyon population, stabilizing anyonic dipoles at temperatures below the charge gap. Furthermore, we investigate neutral anyons in even-denominator FQHE states, which are likely described by non-Abelian wavefunctions. These findings open the door to exploring non-Abelian statistics in neutral anyons, with the potential to reshape future research in topological quantum phases.
@@
Figure 1: Exciton pairing between a pair of Laughlin states. (a) Schematic of the double-layer graphene device used in this geometry to stabilize spatially indirect exciton pairing between quasiparticles and quasiholes across two graphene layers. (b) Schematic of the edgeless Corbino geometry used for transport measurements in parallel flow (PF), counterflow, and drag configurations. (c) Schematic diagram showing FQHE within the CF model, where two flux attachments are included in the CF construction. Blue and red solid lines denote integer values for $\nu^{\ast}_{1}$ and $\nu^{\ast}_{2}$, respectively  Li2019pairing ; Liu2019interlayer . At the intersection of these lines, both layers simultaneously develop a robust charge gap, leading to an incompressible state across the quantum Hall bilayer. (d-e) Transport response along $\nu_{1}^{\ast}=1$, at (d) $B=12$ T and (e) $B=28$ T. The top, middle, and bottom panels show PF conductance $G_{\text{PF}}$, counterflow conductance $G_{\text{CF}}$, and drag ratio $I_{\text{drag}}/I_{\text{drive}}$, respectively. Inset of each panel shows schematic measurement configurations in the parallel flow, counterflow, and drag geometries. Blue, green, and red circles mark different ground state orders.
Figure 2: Pairing between quasiparticle excitations. (a) Arrhenius plot of bulk conductance measured in the parallel flow $G_{\text{PF}}$ (black) and counterflow $G_{\text{CF}}$ (blue) configurations at $\nu_{1}=\nu_{2}=1/3$ and $B$ = 28 T. Inset shows thermally excitated quasiparticles and quasiholes across the charge gap of the Laughlin state. (b) $B$-dependence of counterflow conductance $G_{\text{CF}}$ in logarithmic scale measured at $\nu_{1}=\nu_{2}=1/3$ (Black circles), $2/5$ (blue circles), and $3/7$ (Red circles).
Figure 3: Doping neutral anyons into the Laughlin state. (a) Controlling the population of neutral anyons through layer-asymmetric doping, which is achieved by varying Landau level fillings of layer 1 and 2 by ±$\pm\delta$. This generates equal number of quasiparticles and quasiholes across two graphene layers. (b) Parallel flow conductance $G_{\text{PF}}$ (left), drag ratio $I_{\text{drag}}/I_{\text{drive}}$ (middle), and schematic diagram (right) summarizing the transport response across the $\nu_{\text{total}}$-$\Delta\nu$ map around the Laughlin state at $\nu_{1}=\nu_{2}=1/3$ and $B=28$ T. Near $\Delta\nu$=0, $I_{\text{drag}}$ and $I_{\text{drive}}$ both vanish, resulting in ill-defined drag ratio masked by gray-shaded areas. Black arrows in the right panel mark the extend of the robust charge gap along the direction of $\Delta\nu$ and $\nu_{\text{total}}$. (c) $G_{\text{PF}}$, $G_{\text{CF}}$ (top) and $I_{\text{drag}}/I_{\text{drive}}$ (bottom) as a function of $\Delta\nu$ measured at $\nu_{\text{total}}=2/3$. (d) Arrhenius plot of $G_{\text{PF}}$ (black) and $G_{\text{CF}}$ (blue) measured at $\Delta\nu=0.06$. (e) $G_{\text{PF}}$, $G_{\text{CF}}$ as a function of varying $\nu_{1}$ while $\nu_{2}$ is fixed at $1/3$.
Figure 4: Non-Abelian neutral mode at $\nu_{1}=\nu_{2}=3/8$. (a) A detailed view of the $\nu_{\text{total}}$-$\Delta\nu$ phase space, where black and gray lines indicate CF states with two and four flux attachments, respectively. (b) Parallel flow conductance $G_{\text{PF}}$, (c) counterflow conductance $G_{\text{CF}}$, and (d) drag ratio $I_{\text{drag}}/I_{\text{drive}}$ as a function of varying $\nu_{\text{total}}$ measured at $B=31$ T. Vertical arrows highlight different types of charge-neutral anyons emerging at locations marked by red, blue, and green circles in panel (a). (e) $G_{\text{PF}}$ as a function of $\nu_{\text{total}}$ and $B$, revealing a series of $B$-induced transitions in the FQHE states, marked by horizontal black solid lines. (f) $G_{\text{PF}}$ measured across a portion of the phase space encircled by the red dashed line in panel (a). Vanishing $G_{\text{PF}}$ defines incompressible states marked by black and gray circles in panel (a).
Quantum statistics underpins our understanding of various quantum phases of matter, a cornerstone of condensed matter research. In an electronic system, fermionic statistics are key to describing various quantum phases emerging from a low-temperature electronic systems  Luttinger1960fermi . In contrast, excitons, which are bound states of electrons and holes, provide a pathway to explore bosonic ground states in solid state systems, such as Bose-Einstein condensation  Blatt1962exciton ; Lozovik1975exciton ; Lozovik1976exciton . An even more exotic class of quantum particles, known as anyons, emerges in the context of the fractional quantum Hall effect  Tsui1982FQHE ; Wilczek1982anyon ; Wilczek1982fractional ; Arovas1984fractional ; review-FH ; Picciotto1997 ; Reznikov1999 ; Dolev2008 . Driven by the rich interplay between electron correlation and topology within a partially occupied Landau level (LL), quasiparticle excitations are shown to carry a fractional multiple of the electron’s charge. As a defining characteristic of the FQHE, this fractional charge plays a pivotal role in shaping its transport properties, including the hallmark signature of quantized Hall conductance at fractional values of $e^{2}/h$  Tsui1982FQHE .
Although fractional statistics are often associated with charge-carrying particles  Wilczek1982anyon ; Wilczek1982fractional ; Arovas1984fractional ; Picciotto1997 ; Reznikov1999 ; Dolev2008 , neutral particles can also exhibit anyonic behavior. A unique class of neutral anyons has been discussed in the context of non-Abelian wavefunctions. A notable example is the chiral Majorana mode, which obeys non-Abelian Ising statistics and enables the braiding operations in distinct fusion sectors, a critical feature for topological qubits  Nayak.08 . Recent experiments have leveraged the thermal conductance of such neutral anyons as a key fingerprint to distinguish between candidate topological orders, such as the particle-hole Pfaffian (PH-Pf) and anti-Pfaffian (A-Pf) states  Banerjee2017a ; Banerjee2017b . Unlike quantum phases governed by Coulomb interactions between charged particles, excitonic systems follow fundamentally different dynamics due to their charge neutrality, opening new avenues for exploring exotic quantum states. The charge neutrality also decouples neutral anyons from environmental interference, enhancing their robustness. The experimental realization of neutral anyons, therefore, holds great potential to broaden our understanding of quantum phases of matter with exotic topology, presenting a new frontier for condensed matter physics.
Charge-neutral anyons, by definition, do not carry net charge, rendering conventional charge transport techniques ineffective in probing their behavior. This limitation has driven the development of new methods to study these exotic particles. We report a novel approach for stabilizing, identifying, and controlling charge-neutral anyons in quantum Hall bilayer systems, achieved through excitonic pairing between bulk quasiparticles and quasiholes, both of which exhibit anyonic statistics. This unlocks the intriguing possibility of an anyonic exciton. Using field-effect doping, we precisely control the population of these neutral anyons by introducing equal numbers of bulk quasiparticles and quasiholes into two graphene layers. This induces a transition from a state, devoid of bilayer excitons, to a distinct phase where stable excitonic anyons persist even at temperatures significantly below the charge gap. Our findings highlight a novel ground state consisting of interlayer “anyon dipoles” coexisting with the FQHE in both layers.
Amongst a multitude of neutral anyons observed, we particularly highlight those associated with an even-denominator state, where the likely ground state order is non-Abelian in nature. This leads to neutral excitations obeying non-Abelian statistics. Our findings open new avenues for exploring charge-neutral anyon with exotic properties, providing new insights into the intricate interplay between correlation, topology, and charge fractionalization.
The quantum Hall bilayer system confines charge carriers in closely separated but electrically isolated graphene layers, as illustrated in Fig. 1a  Li2017superfluid ; Liu2017superfluid ; Li2019pairing ; Liu2022crossover . To probe bulk quasiparticles and quasiholes, we pattern the bilayer system into the edgeless Corbino geometry  Nandi2012exciton ; Eisenstein2014 ; Li2019pairing ; Zeng2023solid ; Zhang2024fractionalexciton , which enables three distinct electrical transport configurations (Fig. 1b, also see Methods Section for more detailed discussion). The bulk conductance measurement in the parallel flow (PF) geometry probes the charge gap of the bilayer system, while the counterflow and drag geometries serve as direct probes of charge neutral-modes with non-zero dipole moment, such as interlayer excitons.
@@
When subjected to an out-of-plane magnetic field, the LL filling of each graphene layer, $\nu_{1}$ and $\nu_{2}$, can be independently controlled by adjusting the voltage bias on the top and bottom gate electrodes, $V_{1}$ and $V_{2}$. This configuration allows access to a low-temperature phase space defined by the total filling $\nu_{\text{total}}=\nu_{1}+\nu_{2}$ and the layer imbalance $\Delta\nu=\nu_{1}-\nu_{2}$. We focus on a region of the phase space dominated by the conventional Jain sequence of FQHE states, whose locations are indicated by black circles in Fig. 1c. At these locations, layer 1 is occupied by a Laughlin state at $\nu_{1}=1/3$, while layer 2 is tuned across a sequence of Jain states characterized by $\nu_{2}^{\ast}=\nu_{2}/(1-2\nu_{2})=N$, where $N\in\mathbb{Z}$ defines the sequence. In this work, we use $\nu_{i}^{\ast}$ to denote composite fermion (CF) effective filling $\nu^{\ast}$ in layer $i=1,2$  Jain.03 ; Eisenstein1990FQHE . These CFs are constructed with intralayer flux attachments only, distinguishing them from previous reports of two-component CFs  Li2019pairing ; Zhang2024fractionalexciton .
@@
Fig. 1d plots transport measurement across this sequence of Jain states at $B=12$ T. As shown in the top panel, the PF geometry reveals vanishing bulk conductance, confirming the presence of robust charge gaps at the expected positions of Jain states. Interestingly, the insulating states in the PF geometry exhibits non-zero conductance in the counterflow measurement (middle panel of Fig. 1d), which also coincide with the perfect drag response (bottom panel of Fig. 1d). As a current bias is applied to the drive layer (layer 1), a counter-propagating current of equal magnitude is detected in the drag layer (layer 2), giving rise to a unity drag ratio $I_{\text{drag}}/I_{\text{drive}}=1$. These observations suggest the presence of neutral modes, likely arising from interlayer exciton pairing.
@@
It is important to note that the Jain sequence of FQHE states, characterized by its hierarchical behavior shown in Fig. 1d, is described by well-defined wavefunctions localized within each layer. The absence of interlayer correlation in the CF construction raises an intriguing conundrum regarding the origin of the neutral mode. An interesting $B$-dependence offers an important clue. As shown in Fig. 1e, the neutral mode between a pair of Laughlin state at $\nu_{1}=\nu_{2}=1/3$ vanishes at $B=28$ T. At this large magnetic field, vanishing conductance is observed in both the PF and counterflow geometries (top and middle panels of Fig. 1e). In the drag geometry, both the drive and drag currents approach zero, leading to an ill-defined drag ratio, which is masked with the gray-shaded region in the bottom panel of Fig. 1e.
@@
We propose that the neutral mode arises from exciton pairing between negatively charged quasiparticle in one layer and a positively charged quasihole in the other. As shown in the inset of Fig. 2a, quasiparticles and quasiholes are thermally generated across the charge gap. As such, their population is governed by the temperature in relation to the charge gap. In systems with a large charge gap, the neutral modes are suppressed at low temperatures, but are expected to re-emerge as the temperature increases. This hypothesis is validated by the temperature dependence of $G_{\text{PF}}$ and $G_{\text{CF}}$, measured at $B=28$ T and $\nu_{1}=\nu_{2}=1/3$, as shown in Fig. 2a. The activated behavior of $G_{\text{PF}}$ is directly linked to the charge gap, $\Delta_{\text{charge}}$, which represents the energy required to create thermally excited quasiparticles and quasiholes. Interestingly, the activated behavior of $G_{\text{CF}}$ reveals the energy required to generate a charge-neutral excitation, $\Delta_{\text{exciton}}$, which closely aligns with $\Delta_{\text{charge}}$. This provides direct evidence that both quasiparticles and the neutral mode are generated through the same thermal process.
@@
It should be noted that the Jain states follow a hallmark hierarchical behavior. As the effective magnetic field decreases at larger effective filling, the charge gap is reduced, making it more preferable to generate quasiparticles and quasiholes. This has a direct impact on the transport response in the counterflow channel. As shown in the bottom panel of Fig. 1e, while the Laughlin state exhibits zero counterflow conductance at $B=28$ T, a perfect drag response is recovered as layer 2 is tuned to Jain states at larger effective fillings.
@@
Along the same vein, Fig. 2b plots counterflow conductance $G_{\text{CF}}$ as a function of $B$ for three pairs of Jain states. The presence of charge-neutral modes, indicated by a non-zero $G_{\text{CF}}$, disappears at $\nu_{1}$ = $\nu_{2}$ = 1/3 after $B=15$ T, while persisting up to $B$ = 30 T at $\nu_{1}$ = $\nu_{2}$ = 2/5 and even higher $B$ at $\nu_{1}$ = $\nu_{2}=3/7$. This comparison establishes a general behavior: at each filling, the neutral mode is suppressed as the charge gap is enhanced with increasing $B$, while at each $B$, the neutral mode is enhanced as the charge gap is suppressed at larger effective filling. The direct correlation between the charge gap and neutral mode stability strongly supports the proposed pairing mechanism between quasiparticles.
Together, our observations provide strong evidence for neutral anyons composed of fractionally charged components. The anyonic nature stems from the fact that both the quasiparticles and quasiholes in the Laughlin state share the same braiding phase of $32\pi/$ 3. Consequently, the combined neutral exciton exhibits a braiding phase of $34\pi/$ 3. We note that the quantum statistics of the neutral mode can be tuned by selecting its parent Jain state. For instance, the neutral mode derived from $\nu_{1}=1/3$ and $\nu_{2}=2/3$ is expected to be bosonic, as the quasiparticles and quasiholes exhibit opposite braiding phase of ±$\pm 2\pi/3$ (see Fig. M1 and Method Section II for further details). This results in a robust ground state described by the Bose-Einstein condensation, which is evidenced by the perfect drag response observed at low temperature and $B=28$ T (Fig. 1e).
We note that the $B$-dependence of the neutral mode cannot be accounted for by the effective interlayer separation, $d/\ell_{B}$, a parameter typically used to describe the stability of the exciton condensate at a total filling of one  Eisenstein2014 ; Li2017superfluid ; Liu2017superfluid ; Liu2022crossover . Here, the magnetic length $\ell_{B}$ governs intralayer correlations, while the interlayer separation $d$ defines interlayer correlations. It is important to distinguish between the neutral anyons in Laughlin states, where the charge gap emerges solely from intralayer correlations, and the interlayer excitons of the (111) state Halperin1983 ; Wen1992 , a two-component ground state where interlayer coupling is essential. As a result, $d/\ell_{B}$ is inadequate for describing the properties of neutral anyons.
When quasiparticles are thermally generated, their corresponding neutral mode also behaves as a thermal excitation that disappears at low temperatures, as shown in Fig. 1d and Fig. 2a. In comparison, the presence of neutral anyons as a ground state feature would be highly unusual, potentially leading to exotic phases such as the anyonic superconductor  chen1989anyon ; lee1989anyon . In this work, we show that the low-temperature population of neutral anyons can be controlled through field-effect doping.
@@
As illustrated in Fig. 3a, an equal population of quasiparticles and quasiholes can be generated across the two graphene layers by doping the bilayer to $\nu_{1}=1/3+\delta$ and $\nu_{2}=1/3-\delta$. As quasiparticles and quasiholes pair to form charge-neutral anyons, the charge gap remains robust against this form of doping, which describes a vertical line in phase space as a function of varying layer imbalance $\Delta\nu$, while keeping the total filling fraction $\nu_{\text{total}}$ constant.
@@
The presence of doping-induced excitons is demonstrated by the transport properties around the Laughlin state at $\nu_{1}=\nu_{2}=1/3$. Vanishing parallel flow conductance with $G_{\text{PF}}=0$ highlights a region with a fully developed charge gap, represented by dark blue in the left panel of Fig. 3b. While this region spans a range of $0.20$ in the direction of $\Delta\nu$, it covers a much narrower range of only $0.02$ in the direction of $\nu_{\text{total}}$. This directional dependence aligns well with the unique doping requirement for exciton formation. Tuning along other directions, such as varying $\nu_{\text{total}}$ while holding $\Delta\nu$ fixed, leads to mismatched quasiparticle and quasihole populations, resulting in unpaired charges and a lack of drag response.

### v2

# Bilayer Excitons in the Laughlin Fractional Quantum Hall State

@@
Bilayer Excitons in the Laughlin Fractional Quantum Hall State
@@
Ron Q. Nguyen1    Naiyuan J. Zhang1    Navketan Khurana-Batra1,2    Xiaoxue Liu3    Kenji Watanabe4    Takashi Taniguchi5    D. E. Feldman1,2    J.I.A. Li1 1Department of Physics, Brown University, Providence, Rhode Island 02912, USA 2Brown Theoretical Physics Center, Brown University, Providence, Rhode Island 02912, USA 3Tsung-Dao Lee Institute and School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China 4Research Center for Electronic and Optical Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 5Research Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(July 5, 2025)
The Laughlin state embodies a universal class of fractional quantum Hall effects arising in two-dimensional electron systems subjected to strong perpendicular magnetic fields. Conventionally described by a single-component wavefunction, the Laughlin state features fractionally charged quasiparticles arising from correlations within one electron species. Here, we explore a novel physical situation by introducing inter-species Coulomb coupling between two intra-species Laughlin states in a quantum Hall graphene bilayer structure. Although quasiparticle excitations typically exhibit charge gaps of tens of Kelvin, we observe that this energy scale is significantly lowered through interlayer excitonic pairing between quasiparticles and quasiholes. Identified via transport measurements, these excitons belong to an unprecedented category of charge-neutral anyons, opening a new avenue for investigating exotic quantum statistics and phases of matter.
@@
Figure 1: Exciton pairing between a pair of Laughlin states. (a) Schematic of the Laughlin state in a monolayer, stabilized by intralayer Coulomb interaction. At $1/3$ LL filling, quasiparticle excitations of the Laughlin state carry $1/3$ of electronic charge. (b) Schematic of two Laughlin states in a quantum Hall bilayer, each stabilized by intralayer Coulomb interaction. In this situation, quasiparticles and quasiholes interact through weak interlayer correlation. (c) Schematic diagram showing FQHE within the CF model, where the attachment of two flux quanta is included in the CF construction  Jain.03 . Blue and red solid lines denote integer values for $\nu^{\ast}_{1}$ and $\nu^{\ast}_{2}$, respectively  Li2019pairing ; Liu2019interlayer . At the intersection of these lines, both layers simultaneously develop a robust charge gap, leading to an incompressible state across the quantum Hall bilayer. (d-e) Transport response along $\nu_{1}^{\ast}=1$, at (d) $B=12$ T and (e) $B=28$ T. The top, middle, and bottom panels show PF conductance $G_{\text{PF}}$, counterflow conductance $G_{\text{counter}}$, and drag ratio $I_{\text{drag}}/I_{\text{drive}}$, respectively. Inset of each panel shows schematic measurement configurations in the parallel flow, counterflow, and drag geometries. Blue, green, and red circles mark different ground state orders.
Figure 2: Pairing between quasiparticle excitations. (a,b) Schematic diagrams illustrating (a) thermally excited quasiparticles and quasiholes across the charge gap of the Laughlin state, and (b) suppressed quasiparticle excitation when the charge gap greatly exceeds $k_{B}T$. (c,d) Temperature dependence of (c) drive and drag currents, $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$, in a drag measurement, and (d) the drag ratio $I_{\textrm{drag}}/I_{\textrm{drive}}$. Measurements are performed at $\nu_{1}=\nu_{2}=1/3$ and $B=28$ T. The gray shaded region indicates the low-temperature regime where both currents vanish. The red shaded region marks the onset of simultaneous $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$, yielding a perfect drag ratio. At higher temperatures, $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$ bifurcate, and the drag ratio deviates from unity. (e) Magnetic field dependence of the counterflow conductance $G_{\text{counter}}$ (log scale) measured at $\nu_{1}=\nu_{2}=1/3$ (black circles), $2/5$ (blue circles), and $3/7$ (red circles).
Quantum statistics form the cornerstone for understanding diverse quantum phases of matter in condensed matter physics. Fermionic statistics fundamentally describe electronic systems at low temperatures, underpinning the physics of Fermi liquids Luttinger1960fermi ; Nozieres2018quantumliquids and the quantum Hall effect Halperin2020fractional . When fermions pair, they form composite particles which are bosons, enabling the emergence of low-temperature ground states characterized by Bose–Einstein condensation. Superconductivity, for instance, results from electrons near Fermi surface undergoing Cooper pairing instabilities Cooper2011BCS . Alternatively, excitonic superfluids represent a distinct class of bosonic condensate, resulting from Coulomb-driven electron-hole pairing, establishing an intriguing phenomenon that is actively pursued in condensed matter research Blatt1962exciton ; Lozovik1975exciton ; Lozovik1976exciton .
While excitons are accessible in a wide range of materials, quantum Hall bilayers represent a paradigmatic solid-state platform for stabilizing excitonic condensates. In these bilayer structures, charge carriers reside in two closely spaced but electrically isolated two-dimensional layers. Coulomb attraction across the layers drives electron-hole pairing, stabilizing various emergent excitonic phases extensively studied both experimentally and theoretically Nandi2012exciton ; Eisenstein2014 ; Li2017superfluid ; Liu2017superfluid ; Li2019pairing ; Liu2022crossover ; Eisenstein2004BEC ; Kellogg2005 ; Zeng2023solid ; Zhang2025fractionalexciton ; Kellogg2003 ; Kellogg2004 ; Tutuc2004counterflow . Among the most extensively examined phases, an excitonic condensate emerges at a total Landau level (LL) filling factor of one, accurately described by the (111) wavefunction Halperin1983 ; Wen1992 , representing an excitonic analogue of the integer quantum Hall effect state.
More recently, it was discovered that interlayer excitons can coexist with fractional quantum Hall states, indicating that excitonic phenomena are more ubiquitous within the quantum Hall regime than previously recognized. In the regime with strong interlayer correlation, fractional quantum Hall states can be described by two-component wavefunctions, notably the (333) state Halperin1983 ; Wen1992 , serving as the fractional analogue to the (111) state. Here, interlayer excitons naturally emerge as essential constituents of the ground-state construction Zhang2025fractionalexciton .
In this work, we report a new pathway for creating interlayer excitons by introducing Coulomb coupling between two Laughlin states. While the Laughlin state is traditionally regarded as a single-component state, we show that Coulomb coupling between two such states generates interlayer excitons at energies substantially lower than the charge gap. As shown in Fig. 1a, quasiparticle excitation of the Laughlin state carries $1/3$ of an electron charge and obeys anyonic quantum statistics Tsui1982FQHE ; halperin1984 ; Arovas1984fractional ; review-FH ; Picciotto1997 ; Saminadayar1997 ; Reznikov1999 ; Bartolomei2020fractional ; Nakamura2020Anyon ; texp1 ; texp2 ; Willett2023interference . We propose that interlayer excitons of two Laughlin states results from pairing between fractionally charged quasiparticles and quasiholes (Fig. 1b).
@@
The phase space of a quantum Hall bilayer is defined by the Landau level fillings of each graphene layer, $\nu_{1}$ and $\nu_{2}$, which can be independently controlled by adjusting the voltage bias on the top and bottom gate electrodes. Figure 1c presents a schematic phase diagram, where the axes are redefined by the total filling $\nu_{\text{total}}=\nu_{1}+\nu_{2}$ and the layer imbalance $\Delta\nu=\nu_{1}-\nu_{2}$. We focus on the weak-coupling regime dominated by the conventional Jain sequence of FQHE states. Here, red and blue solid lines indicate the expected locations of single-component FQHE states in layers 1 and 2, respectively. Black circles mark a representative set of intersections, where layer 1 is tuned to a Laughlin state at $\nu_{1}=1/3$ and layer 2 is swept across a sequence of Jain states characterized by the effective composite fermion (CF) filling $\nu_{2}^{\ast}=\nu_{2}/(1-2\nu_{2})=N$ with $N\in\mathbb{Z}$ Jain.03 ; Eisenstein1990FQHE . These CFs are constructed using intralayer flux attachment only, distinguishing them from previous reports of two-component FQHE Li2019pairing ; Zhang2025fractionalexciton . In the following, we use $\nu_{i}^{\ast}$ to denote the CF effective filling in layer $i=1,2$.
@@
Across the Jain sequence highlighted in Fig. 1c, the presence of interlayer excitons is revealed through their distinctive transport responses. We employ three distinct electrical transport configurations in bilayer samples patterned into an edgeless Corbino geometry (see Methods for details) Nandi2012exciton ; Eisenstein2014 ; Li2019pairing ; Zeng2023solid ; Zhang2025fractionalexciton . At $B=12$ T, the parallel-flow (PF) geometry exhibits vanishing bulk conductance at $\nu_{2}^{\ast}=N$ (top panel of Fig. 1d), confirming robust charge gaps associated with Jain states. Remarkably, these insulating states in the PF geometry exhibit non-zero conductance in the counterflow configuration (middle panel of Fig. 1d), coinciding with perfect drag responses (bottom panel of Fig. 1d). When a current bias is applied to the drive layer (layer 1), a counter-propagating current of equal magnitude is detected in the drag layer (layer 2), resulting in a unity drag ratio $I_{\text{drag}}/I_{\text{drive}}=1$. This combination of transport signatures is in excellent agreement with prior observations of the (111) state, offering unambiguous evidence for the existence of interlayer excitons across the Jain sequence.
@@
Interestingly, signatures of exciton pairing between a pair of Laughlin state vanishes at $B=28$ T. As shown in Fig. 1e, vanishing conductance is observed in both the PF and counterflow geometries at $\nu_{1}=\nu_{2}=1/3$ (top and middle panels of Fig. 1e). In the drag geometry, both the drive and drag currents approach zero, leading to an ill-defined drag ratio, which is masked with the gray-shaded region in the bottom panel of Fig. 1e.
The Jain sequence of FQHE states, characterized by its hierarchical behavior shown in Fig. 1d-e, is described by well-defined wavefunctions localized within each layer. The absence of interlayer correlation in the ground-state wavefunction raises an intriguing conundrum regarding the origin of interlayer excitons.
@@
The magnetic field dependence of the transport response, as illustrated in Fig. 1d-e, provides an important clue. We propose that exciton pairing occurs between a negatively charged quasiparticle in one layer and a positively charged quasihole in the other. As shown in the inset of Fig. 2a, such quasiparticles and quasiholes are thermally generated across the charge gap. Consequently, their population is governed by the temperature relative to the charge gap. In systems with a large charge gap, such as the situation at $B=28$ T, these excitons are suppressed at low temperatures but are expected to re-emerge as the gap diminishes or as the temperature increases.
This hypothesis is validated by the temperature dependence of the drag response. At $B=28$ T and $\nu_{1}=\nu_{2}=1/3$, drag measurement produces zero $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$ in the low-temperature region, as shown in Fig. 2c. This is consistent with the absence of excitons, and both layers behave like insulators in the bulk. With increasing temperature, quasiparticle excitations are generated across the charge gap, evidenced by non-zero current in both drive and drag layers. Remarkably, the onset of current across two layers recovers a perfect drag ratio, as shown in the red shaded area in Fig. 2d. This observation offers strong support for exciton pairing from thermally excited quasiparticle and quasihole. Upon further increasing $T$, $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$ bifurcate and the drag ratio start to deviate from $1$, indicating that exciton pairing is suppressed at high temperature.
Since the transport response in the parallel flow configuration does not couple to interlayer excitons, the activated behavior in $G_{\text{PF}}$ offers a direct measurement of the underlying charge gap. As illustrated by Fig. S1, the charge gap associated with the Laughlin state at $B=28$T is around $\Delta=40.5$ K, substantially higher than the temperature of current onset in the drag measurement. This discrepancy suggests that exciton pairing lowers the energy cost for creating quasiparticle excitations.
@@
Moreover, the Jain states follow a hallmark hierarchical behavior. As the effective magnetic field decreases at larger effective filling, the charge gap is reduced, making it more preferable to generate quasiparticles and quasiholes. This has a direct impact on the transport response in the counterflow channel. As shown in the bottom panel of Fig. 1e, while the Laughlin state exhibits zero counterflow conductance at $B=28$ T, a perfect drag response is recovered as layer 2 is tuned to Jain states at larger effective fillings.
@@
Along the same vein, Fig. 2e plots counterflow conductance $G_{\text{counter}}$ as a function of $B$ measured at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=1$, $2$, and $3$. At $\nu_{1}^{\ast}=\nu_{2}^{\ast}=1$, exciton population disappears after $B=15$ T, evidenced by vanishing $G_{\text{counter}}$. On the other hand, exciton pairing persists up to $B$ = 30 T at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=2$ and even higher $B$ at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=3$. This comparison establishes a general behavior: at each filling, exciton pairing is suppressed as the charge gap is enhanced with increasing $B$, while at each $B$, excitons become more stable as the charge gap is suppressed at larger effective filling. Combined, our observations strongly support the proposed pairing mechanism between quasiparticles.
Figure 3: Doping neutral anyons into the Laughlin state. (a) Controlling the population of neutral anyons through layer-asymmetric doping, which is achieved by varying Landau level fillings of layer 1 and 2 by ±$\pm\delta$. This generates equal number of quasiparticles and quasiholes across two graphene layers. (b) Parallel flow conductance $G_{\text{PF}}$ (left), drag ratio $I_{\text{drag}}/I_{\text{drive}}$ (middle), and schematic diagram (right) summarizing the transport response across the $\nu_{\text{total}}$-$\Delta\nu$ map around the Laughlin state at $\nu_{1}=\nu_{2}=1/3$ and $B=28$ T. Near $\Delta\nu$=0, $I_{\text{drag}}$ and $I_{\text{drive}}$ both vanish, resulting in ill-defined drag ratio masked by gray-shaded areas. Black arrows in the right panel mark the extend of the robust charge gap along the direction of $\Delta\nu$ and $\nu_{\text{total}}$. (c) $G_{\text{PF}}$, $G_{\text{counter}}$ (top) and $I_{\text{drag}}/I_{\text{drive}}$ (bottom) as a function of $\Delta\nu$ measured at $\nu_{\text{total}}=2/3$. (d) Arrhenius plot of $G_{\text{PF}}$ (black) and $G_{\text{counter}}$ (blue) measured at $\Delta\nu=0.06$. (e) $G_{\text{PF}}$, $G_{\text{counter}}$ as a function of varying $\nu_{1}$ while $\nu_{2}$ is fixed at $1/3$.
Figure 4: Non-Abelian neutral mode at $\nu_{1}=\nu_{2}=3/8$. (a) A detailed view of the $\nu_{\text{total}}$-$\Delta\nu$ phase space, where black and gray lines indicate CF states with two and four flux attachments, respectively. (b) Parallel flow conductance $G_{\text{PF}}$, (c) counterflow conductance $G_{\text{counter}}$, and (d) drag ratio $I_{\text{drag}}/I_{\text{drive}}$ as a function of varying $\nu_{\text{total}}$ measured at $B=31$ T. Vertical arrows highlight different types of charge-neutral anyons emerging at locations marked by red, blue, and green circles in panel (a). (e) $G_{\text{PF}}$ as a function of $\nu_{\text{total}}$ and $B$, revealing a series of $B$-induced transitions in the FQHE states, marked by horizontal black solid lines. (f) $G_{\text{PF}}$ measured across a portion of the phase space encircled by the red dashed line in panel (a). Vanishing $G_{\text{PF}}$ defines incompressible states marked by black and gray circles in panel (a).
In the above mentioned scenario, quasiparticles, quasiholes, and the resulting excitons are all thermal excitations. In the following, we show that such excitons can also be generated as a feature of the low-temperature ground state, which is realized through independently control field-effect doping across two graphene layers.
@@
As illustrated in Fig. 3a, an equal population of quasiparticles and quasiholes can be generated across the two graphene layers by doping the bilayer to $\nu_{1}=1/3+\delta$ and $\nu_{2}=1/3-\delta$. This defines a vertical line in phase space at constant total filling fraction $\nu_{\text{total}}=2/3$, where $\delta$ is directly linked to layer imbalance $\Delta\nu$.
@@
The presence of doping-induced excitons is demonstrated by the transport properties around $\nu_{1}=\nu_{2}=1/3$. Vanishing parallel flow conductance with $G_{\text{PF}}=0$ highlights a region with a fully developed charge gap, represented by dark blue in the left panel of Fig. 3b. While this region spans a range of $0.20$ in the direction of $\Delta\nu$, it covers a much narrower range of only $0.02$ in the direction of $\nu_{\text{total}}$. This directional dependence aligns well with the unique doping requirement for exciton formation. Tuning along other directions, such as varying $\nu_{\text{total}}$ while holding $\Delta\nu$ fixed, leads to mismatched quasiparticle and quasihole populations, resulting in unpaired charges and a lack of drag response.